OpenStudy (anonymous):

If z = sqrt2 - 2i, then |z| = _____.

OpenStudy (anonymous):

sqrt 2 sqrt 6 3 6

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

give me a sec my computer a little slow,

OpenStudy (anonymous):

to be honest I dont really know but I think it is 3 just from what i see

jimthompson5910 (jim_thompson5910):

They gave you \(\Large z = \sqrt{2} - 2i\) right?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

so a = sqrt(2) ---> a^2 = 2 b = -2 ---> b^2 = 4 |z| = sqrt( a^2 + b^2 ) |z| = sqrt(2 + 4) |z| = sqrt( 6 )